OpenStudy (toxicsugar22):

Caffenine leaves the body at a rate of 15.6% each hour. Your first cup of coffee in the morning has 100 mg of caffenine. How long will it take before you have 50 mg and 25 mg of that caffenine in your body.

OpenStudy (anonymous):

An equation to model how much caffeine you have left in your body would be \[c=100*(.844)^{x}\] So set c=50 and then 25. \[50=100*(.844)^{x}\] x=4.08689 hours \[25=100*(.844)^{x}\] x=8.17377

OpenStudy (toxicsugar22):

thanks

OpenStudy (toxicsugar22):

how did get .844

OpenStudy (anonymous):

15.6% is equal to .156, but that's the amount LEAVING your body. The amount LEFT IN your body would be 100%-15.6%, or 1-0.156= .844. That help?

OpenStudy (toxicsugar22):

Yes thanks you so much

OpenStudy (toxicsugar22):

can u help me with another question please

OpenStudy (anonymous):

What is it?

OpenStudy (toxicsugar22):

Dichloro is a pesticide that was used in the middle decades of the 20th century to control maleria. Afterr 1945, it was also widely used on crops in the U, and as much as one ton be sprayed on a single cotton field. Hoowever after the toxic, effects of Dichloro on the evvironment began to appear, the chemical was banned in 1972. A common estimate for the half-life in the soil is 15 years. Write a decay law for Dichloro in the soil and use A for the initial concentration of Dicloro, and use t for the input variable

OpenStudy (toxicsugar22):

do u understand

OpenStudy (anonymous):

Not really, I don't remember the rules for half-lives.

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

thanks anyway

OpenStudy (anonymous):

Sorry!